# Bounding field excursions along null geodesics with applications to cosmology

### v1

@@ 5 Quantum effects
where the implicit Newton’s constant is the minimal value permitted within the relevant EFT regime. The kinetic term of the scalar field is also subject to renormalization, affecting the definition of $d(X_{2},X_{1})$. This renormalization should also be considered carefully, perhaps by a similar method to what we proposed for $G_{N}$, using the most conservative effective value of the parameter in the RG flow.
@@
A proof of the SQFC, and therefore the QFEB, is beyond the scope of this paper. Even the validity of the original QFC is still unproven in quantum situations that go beyond the $G\to 0$ (QNEC) limit. [^21] In fact, Ref. [59] suggests that the QFC, and hence the SQFC, may be violated for certain large-$N$ CFTs coupled to gravity for specific choices of spacelike slicing. Nevertheless, it is noteworthy that the QFEB holds in spacetimes that do satisfy the SQFC.
@@
Finally, we comment on the implications of quantum effects for phenomenology. While the QFEB is interesting, it appears to have limited utility in concrete phenomenological applications. Instead, the central question is under what circumstances quantum corrections could invalidate the arguments of Section 4.3. For such an invalidation to occur, a model must satisfy two conditions simultaneously: it must exhibit genuinely “fast” field excursions and feature quantum effects that dominate over the scalar field’s kinetic term in the Raychaudhuri equation. By “fast”, we mean that the FEB is close to saturation, or at least some order one fraction away from saturation. If the FEB is far from being saturated, violating it in a spacetime with a positive cosmological constant would require extraordinarily large quantum corrections, strong enough to generate focusing effects that rival the redshift from cosmic expansion, which seems highly implausible. Conversely, suppose quantum effects do not dominate over the kinetic term. In that case, one still generally obtains an approximate linear bound on the field excursion in terms of the change in e-folds, even if the original FEB is perturbatively violated. [^22] Even when the quantum contribution is comparable to the kinetic term, a linear bound can still persist, though with a modified numerical coefficient. Although we cannot strictly exclude the possibility of models in which both criteria are met, such scenarios would appear to be highly exotic. Notably, the first condition is not obeyed in models obeying the slow-roll approximation, [^23] Although certain cosmic microwave background measurements place strong upper bounds on the slow-roll parameter [60], these constraints apply only within specific models and to the relatively recent cosmological history, roughly the last 60 e-folds of inflation. which encompasses most scenarios of eternal inflation [1, 2, 51]. Of course, by definition, eternal inflation involves arbitrarily large values of $\Delta N$, but inflation is still past-incomplete [61], and it is still interesting to bound the amount of excursion that can take place in some finite number of e-folds.
@@ 6 Discussion
In the context of inflation, our result bears a close resemblance to the Lyth bound [62], which provides a lower limit on the scalar field excursion in terms of the tensor-to-scalar ratio $r$ and the number of e-folds,
@@
When formulating the quantum generalization of the FEB, we were required to assume the strengthened QFC. The strengthened QFC is a novel and intriguing condition on gravitational theories, deserving exploration in its own right. In particular, it would be valuable to investigate explicit examples in higher-dimensional settings to test its validity. Moreover, just as the original QFC led to the QNEC [63, 64, 65, 66] in the non-gravitational limit, it is natural to ask whether the SQFC could imply new, nontrivial constraints on non-gravitational quantum field theories.
@@
can contribute. In the presence of such terms, it is unclear whether a version of the FEB continues to hold. These terms are particularly sensitive to variations of the field in directions transverse to the null geodesic. Interestingly, the Wilson coefficients of such operators are constrained by positivity bounds [67].
One might ask what form a bound based solely on the area would take. Unlike the FEB, a bound expressed purely in terms of the area at the endpoints of a congruence would have the appealing feature of depending only on data defined on the initial and final spacelike slices bounding the light ray. A natural candidate, consistent with known examples [12, 13, 14, 15], is
@@
for some constant $\alpha$. [^24] Eq. (64) is a classical inequality and can be violated in spacetimes that do not satisfy the null energy condition (NEC). This expression can be viewed as a covariant generalization of the proposal in Ref. [14] that field excursions in a localized region of spacetime should be bounded by the volume of the spacetime region. [^25] A bound in terms of spatial volume is excluded for the same reasons that an entropy bound in terms of volume is inconsistent [68, 69]. However, Eq. (64) fails near naked timelike singularities, such as the example discussed in Section 3, at a finite affine distance from the singularity. [^26] We thank Craig Clark for emphasizing this point. Thus, any derivation of Eq. (64) would necessarily rely on, or amount to a proof of, some version of cosmic censorship [70], a notoriously subtle and unresolved issue. [^27] See Ref. [71] for a recent proposal. For example, a possible counterexample to the area bound (64) could be constructed by taking a spacetime with a naked timelike singularity, excising an arbitrarily small region around the singularity, and replacing it with a nonsingular spacetime. However, we show how energy conditions can exclude a large family of these counterexamples in Appendix A.
It would be interesting to explore whether the bounds derived in this paper have any implications for the swampland program. Despite their relation to certain swampland conjectures, our bounds are conceptually distinct. For example, the infinite distance conjecture [72] concerns the behavior of effective field theories near points at infinite distance in the vacuum manifold, whereas the FEB places local constraints on the ability to probe large distances in field space. Similarly, although the trans-Planckian censorship conjecture [73] bears a superficial resemblance to the FEB, it pertains to the suppression of sub-Planckian quantum fluctuations becoming classical, rather than to classical trans-Planckian field excursions in the vacuum manifold.
@@ Acknowledgments
We thank Carolina Figueiredo and Qianshu Lu for their collaboration in the early stages of this work. We thank Chris Akers, Nima Arkani-Hamed, Scott Collier, Åsmund Folkestad, Sergio Hernández-Cuenca, Patrick Jefferson, Zohar Kormagodski, Jonah Kudler-Flam, Juan Maldacena, Jonathan Source, Douglas Stanford, Evita Verheijden, Edward Witten, and Bowen Zhao for helpful discussions. AH is grateful to the Simons Foundation as well as the Edward and Kiyomi Baird Founders’ Circle Member Recognition for their support. Aron Wall was supported by the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime”, the STFC grant ST/P000681/1 “Particles, Fields and Extended Objects”. He was also supported by NSF grant PHY-2207584 during his sabbatical at the IAS.
@@ Appendix A Restrictions on cutting out naked singularities
There is some reason to hope for an area bound on field excursions under the assumption of cosmic censorship. However, one challenge in using cosmic censorship to prove Eq. (64) is that one can naively excise a region near a naked singularity and still find that Eq. (64) is violated. Remarkably, as first suggested in Ref. [11], energy conditions imply that a finite region surrounding the singularity must also be removed when excising the singularity itself. The hope is that this minimal cut-out region captures all the loci where Eq. (64) fails. If so, this would suggest that a proof of Eq. (64) might be possible based on cosmic censorship and local energy conditions. While we cannot rigorously establish this, we provide supporting evidence when the parameter $\beta$ is sufficiently small.
To see why a minimal region must be excised along with the naked singularity, we appeal to the concept of positive quasi-local mass. Suppose we excise a spherical region of radius $r$ centered on the singularity and replace it with another spacetime patch. The quasi-local mass of the inserted region can often be computed from boundary data alone, by Stokes’ theorem, and is therefore independent of the interior geometry. However, defining a positive-definite quasi-local mass in general spacetimes is notoriously subtle. To sidestep this difficulty, we assume the spacetime is stationary and radially symmetric, with non-negative energy density. Under these assumptions, the Hawking quasi-local mass [74, 75, 76] enclosed by a sphere centered at the origin is guaranteed to be positive [77]. If the Hawking mass becomes negative sufficiently close to a naked timelike singularity, it implies that any sewn-in spacetime preserving positive energy conditions must exclude a neighborhood of the singularity.
We define the quasi-local energy using the Hawking quasi-mass instead of the quasi-local mass in Ref. [11], which is not positive-definite away from the weak field regime. Given the metric
@@
the Hawking quasi-local mass is [78]
@@
Again, Eq. (66) is positive definite assuming that the matter-energy density is positive definite [77]. In fact, Ref. [77] shows that the Hawking quasi-local mass obeys the stronger bound,

### v2

@@ 5 Quantum effects
where the implicit Newton’s constant is the minimal value permitted within the relevant EFT regime. The kinetic term of the scalar field is also subject to renormalization, affecting the definition of $d(X_{2},X_{1})$. This renormalization should also be considered carefully, perhaps by a similar method to what we proposed for $G_{\text{N}}$, using the most conservative effective value of the parameter in the RG flow.
@@
A proof of the SQFC, and therefore the QFEB, is beyond the scope of this paper. Even the validity of the original QFC is still unproven in quantum situations that go beyond the $G_{\text{N}}\to 0$ (QNEC) limit. [^21] In fact, Ref. [Shahbazi-Moghaddam:2022hbw] suggests that the QFC, and hence the SQFC, may be violated for certain large-$N$ CFTs coupled to gravity for specific choices of spacelike slicing. Nevertheless, it is noteworthy that the QFEB holds in spacetimes that do satisfy the SQFC.
@@
Finally, we comment on the implications of quantum effects for phenomenology. While the QFEB is interesting, it appears to have limited utility in concrete phenomenological applications. Instead, the central question is under what circumstances quantum corrections could invalidate the arguments of Section 4.3. For such an invalidation to occur, a model must satisfy two conditions simultaneously: it must exhibit genuinely “fast” field excursions and feature quantum effects that dominate over the scalar field’s kinetic term in the Raychaudhuri equation. By “fast”, we mean that the FEB is close to saturation, or at least some order one fraction away from saturation. If the FEB is far from being saturated, violating it in a spacetime with a positive cosmological constant would require extraordinarily large quantum corrections, strong enough to generate focusing effects that rival the redshift from cosmic expansion, which seems highly implausible. Conversely, suppose quantum effects do not dominate over the kinetic term. In that case, one still generally obtains an approximate linear bound on the field excursion in terms of the change in e-folds, even if the original FEB is perturbatively violated. [^22] Even when the quantum contribution is comparable to the kinetic term, a linear bound can still persist, though with a modified numerical coefficient. Although we cannot strictly exclude the possibility of models in which both criteria are met, such scenarios would appear to be highly exotic. Notably, the first condition is not obeyed in models obeying the slow-roll approximation, [^23] Although certain cosmic microwave background measurements place strong upper bounds on the slow-roll parameter [Planck:2018jri], these constraints apply only within specific models and to the relatively recent cosmological history, roughly the last 60 e-folds of inflation. which encompasses most scenarios of eternal inflation [1983veu..conf..251S, Vilenkin:1983xq, Guth:2007ng]. Of course, by definition, eternal inflation involves arbitrarily large values of $\Delta N$, but inflation is still past-incomplete [Borde:2001nh], and it is still interesting to bound the amount of excursion that can take place in some finite number of e-folds.
@@ 6 Discussion
In the context of inflation, our result bears a close resemblance to the Lyth bound [Lyth:1996im], which provides a lower limit on the scalar field excursion in terms of the tensor-to-scalar ratio $r$ and the number of e-folds,
@@
When formulating the quantum generalization of the FEB, we were required to assume the strengthened QFC. The strengthened QFC is a novel and intriguing condition on gravitational theories, deserving exploration in its own right. In particular, it would be valuable to investigate explicit examples in higher-dimensional settings to test its validity. Moreover, just as the original QFC led to the QNEC [Bousso:2015wca, Wall:2017blw, Balakrishnan:2017bjg, Ceyhan:2018zfg] in the non-gravitational limit, it is natural to ask whether the SQFC could imply new, nontrivial constraints on non-gravitational quantum field theories.
@@
can contribute. In the presence of such terms, it is unclear whether a version of the FEB continues to hold. These terms are particularly sensitive to variations of the field in directions transverse to the null geodesic. Interestingly, the Wilson coefficients of such operators are constrained by positivity bounds [Adams:2006sv].
One might ask what form a bound based solely on the area would take. Unlike the FEB, a bound expressed purely in terms of the area at the endpoints of a congruence would have the appealing feature of depending only on data defined on the initial and final spacelike slices bounding the light ray. A natural candidate, consistent with known examples [Dolan:2017vmn, Draper:2019zbb, Draper:2019utz, Delgado:2022dkz], is
@@
for some constant $\alpha$. [^24] Eq. (64) is a classical inequality and can be violated in spacetimes that do not satisfy the null energy condition (NEC). This expression can be viewed as a covariant generalization of the proposal in Ref. [Draper:2019utz] that field excursions in a localized region of spacetime should be bounded by the volume of the spacetime region. [^25] A bound in terms of spatial volume is excluded for the same reasons that an entropy bound in terms of volume is inconsistent [Bousso:1999xy, Bousso:2002ju]. However, Eq. (64) fails near naked timelike singularities, such as the example discussed in Section 3, at a finite affine distance from the singularity. [^26] We thank Craig Clark for emphasizing this point. Thus, any derivation of Eq. (64) would necessarily rely on, or amount to a proof of, some version of cosmic censorship [Penrose:1969pc], a notoriously subtle and unresolved issue. [^27] See Ref. [Engelhardt:2024hpe] for a recent proposal. For example, a possible counterexample to the area bound (64) could be constructed by taking a spacetime with a naked timelike singularity, excising an arbitrarily small region around the singularity, and replacing it with a nonsingular spacetime. However, we show how energy conditions can exclude a large family of these counterexamples in Appendix A.
It would be interesting to explore whether the bounds derived in this paper have any implications for the swampland program. Despite their relation to certain swampland conjectures, our bounds are conceptually distinct. For example, the infinite distance conjecture [Ooguri:2006in] concerns the behavior of effective field theories near points at infinite distance in the vacuum manifold, whereas the FEB places local constraints on the ability to probe large distances in field space. Similarly, although the trans-Planckian censorship conjecture [Bedroya:2019snp] bears a superficial resemblance to the FEB, it pertains to the suppression of sub-Planckian quantum fluctuations becoming classical, rather than to classical trans-Planckian field excursions in the vacuum manifold.
@@ Acknowledgments
We thank Carolina Figueiredo and Qianshu Lu for their collaboration in the early stages of this work. We thank Chris Akers, Nima Arkani-Hamed, Scott Collier, Åsmund Folkestad, Sergio Hernández-Cuenca, Patrick Jefferson, Zohar Kormagodski, Jonah Kudler-Flam, Juan Maldacena, Jonathan Source, Douglas Stanford, Evita Verheijden, Edward Witten, and Bowen Zhao for helpful discussions. AH is grateful to the Simons Foundation as well as the Edward and Kiyomi Baird Founders’ Circle Member Recognition for their support. Aron Wall was supported by the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime”, and the STFC grant ST/X000664/1 “Quantum Fields, Quantum Gravity and Quantum Particles”. He was also supported by NSF grant PHY-2207584 during his sabbatical at the IAS.
@@ Appendix A Restrictions on cutting out naked singularities
There is some reason to hope for an area bound on field excursions under the assumption of cosmic censorship. However, one challenge in using cosmic censorship to prove Eq. (64) is that one can naively excise a region near a naked singularity and still find that Eq. (64) is violated. Remarkably, as first suggested in Ref. [Nicolis:2008wh], energy conditions imply that a finite region surrounding the singularity must also be removed when excising the singularity itself. The hope is that this minimal cut-out region captures all the loci where Eq. (64) fails. If so, this would suggest that a proof of Eq. (64) might be possible based on cosmic censorship and local energy conditions. While we cannot rigorously establish this, we provide supporting evidence when the parameter $\beta$ is sufficiently small.
To see why a minimal region must be excised along with the naked singularity, we appeal to the concept of positive quasi-local mass. Suppose we excise a spherical region of radius $r$ centered on the singularity and replace it with another spacetime patch. The quasi-local mass of the inserted region can often be computed from boundary data alone, by Stokes’ theorem, and is therefore independent of the interior geometry. However, defining a positive-definite quasi-local mass in general spacetimes is notoriously subtle. To sidestep this difficulty, we assume the spacetime is stationary and radially symmetric, with non-negative energy density. Under these assumptions, the Hawking quasi-local mass [Misner:1964je, Hernandez:1966zia, Hawking:1968qt] enclosed by a sphere centered at the origin is guaranteed to be positive [Christodoulou1988SomeRO]. If the Hawking mass becomes negative sufficiently close to a naked timelike singularity, it implies that any sewn-in spacetime preserving positive energy conditions must exclude a neighborhood of the singularity.
We define the quasi-local energy using the Hawking quasi-mass instead of the quasi-local mass in Ref. [Nicolis:2008wh], which is not positive-definite away from the weak field regime. Given the metric
@@
the Hawking quasi-local mass is [Faraoni:2020mdf]
@@
Again, Eq. (66) is positive definite assuming that the matter-energy density is positive definite [Christodoulou1988SomeRO]. In fact, Ref. [Christodoulou1988SomeRO] shows that the Hawking quasi-local mass obeys the stronger bound,